(2.5 multiplied by 10 to the 7th power)+(5.5 multiplied by 10 to the 6th power): (2.5 multiplied by 10 to the 7th power)+(5.5 multiplied by 10 to the 6th power)
Answer by checkley77(3848)

2.5*10^7+5.5*10^6
25,000,000
+5,500,000
-----------------
30,500,000
or 3.05*10^7 answer.
Question 167069: (2.5 multiplied by 10 to the 7th power)+(5.5 multiplied by 10 to the 6th power): (2.5 multiplied by 10 to the 7th power)+(5.5 multiplied by 10 to the 6th power)
Answer by Alan3354(1926)

(2.5 multiplied by 10 to the 7th power)+(5.5 multiplied by 10 to the 6th power)
--------------------
To add or subract, the elements have to have the same exponent. Change the 1st one to 6th power.
2.5*10^7 + 5.5*10^6
= 25*10^6 + 5.5*10^6
= 30.5*10^6
= 3.05*10^7 (standard scientific notation)
Also, 3.05E7 (often used by programmers)